# CodeSense: a Real-World Benchmark and Dataset for Code Semantic Reasoning

### v2

@@ Abstract
Understanding and reasoning about code semantics is essential for enhancing code LLMs’ abilities to solve real-world software engineering (SE) tasks. Although several code reasoning benchmarks exist, most rely on synthetic datasets or educational coding problems and focus on coarse-grained reasoning tasks such as input/output prediction, limiting their effectiveness in evaluating LLMs in practical SE contexts. To bridge this gap, we propose CodeSense , the first benchmark that makes available a spectrum of fine-grained code reasoning tasks concerned with the software engineering of real-world code. We collected Python, C and Java software projects from real-world repositories. We executed tests from these repositories, collected their execution traces, and constructed a ground truth dataset for fine-grained semantic reasoning tasks. We then performed comprehensive evaluations on state-of-the-art LLMs. Our results show a clear performance gap for the models to handle fine-grained reasoning tasks. Although prompting techniques such as chain-of-thought and in-context learning helped, the lack of code semantics in LLMs fundamentally limit models’ capabilities of code reasoning. Besides dataset, benchmark and evaluation, our work produced an execution tracing framework and tool set that make it easy to collect ground truth for fine-grained SE reasoning tasks, offering a strong basis for future benchmark construction and model post training. Our code and data are located at https://codesense-bench.github.io/.
@@ 1 Introduction
Semantic code reasoning—the capacity to understand and predict the behavior of software—is a core requirement underpinning a wide range of complex software engineering (SE) tasks, including test input generation, vulnerability detection, fault localization, bug repair, refactoring, and functional verification. Unlike syntactic pattern matching, which may rely on token-level similarity or statistical regularities, semantic reasoning ("codesense") entails a deep, execution-oriented understanding of how software operates. Although code semantics can be expressed in many ways, in practice, developers engage in semantic reasoning through tasks like predicting a function’s input-output behavior, tracing variable values, analyzing control flow paths, identifying loop invariants, etc. This form of reasoning aligns with formal definitions from programming language theory—particularly operational semantics, which models step-by-step execution, and axiomatic semantics, which uses logical assertions to describe program properties. Such reasoning tasks also reflect the real-world demands placed on developers and provide a natural grounding for their day-to-day work.
@@
To this end, we propose CodeSense, a benchmark for fine-grained code semantic reasoning, constructed from real-world GitHub projects in Python, C, and Java (see Table 1). CodeSense  introduces a spectrum of reasoning tasks at statement, code-block, and function levels, targeting essential semantic properties frequently needed across SE activities. For example, predicting loop iteration counts is critical for input/output prediction, performance analysis, and detecting infinite loops (e.g., denial-of-service vulnerabilities). Branch condition prediction and reasoning about pointers in C code are important for test input generation and memory safety assurance. As illustrated in Figures 1(a) and 1(b), fine-grained semantic reasoning about arithmetic operations, control flow, and API semantics is foundational for a variety of SE applications. Prior work (Ding et al., 2023a; 2024) has shown that incorporating semantic signals during training improves model performance on code generation, branch prediction, code clone detection, program repair and vulnerability detection tasks, motivating our design of CodeSense to comprehensively evaluate models’ capabilities for semantic reasoning. Their experimental results showed that even using dynamic values collected from small implementations (<100 lines of code), the models are able to improve downstream tasks for real-world code. Similarly, Valgrind  (Nethercote and Seward, 2007), a widely used tool and deployed for more than 20 years (Nethercote, 2022), uses dynamic values for reasoning tasks, such as detecting and debugging memory errors, and optimizing programs.
@@
Table 1: Optimal design space of code reasoning benchmarks, ( ○ denotes not support, ◐ denotes partial support, and ● denotes fully support)
Benchmark Real-world Projects Real-world Tests Multilingual Fine-grained Semantic Reasoning
CruxEval
CruxEval-X
REval
CodeMind
@@
Using our benchmark, we evaluated 14 state-of-the-art (SOTA) LLMs and investigated six research questions regarding the models’ code semantics reasoning capabilities. Previous work has shown that models did not perform well on code reasoning tasks such as input/output prediction (Gu et al., 2024) and vulnerability detection (Steenhoek et al., 2025); to understand why models fail and identify places for improvement, we investigate: RQ1: Does increasing code size make semantic reasoning more difficult? RQ2: Which types of program statements are easier or harder for models to reason about? RQ3: How do models perform on code properties critical for SE tasks, such as predicting pointer aliasing, loop iteration counts, and branch conditions? RQ4: How effective are different prompting strategies in improving semantic reasoning? RQ5: Can models reason approximately when exact values or semantics are hard to infer? RQ6: Do models perform better on some programming languages than others?
Our results reveal that current LLMs, including SOTA models like Claude 3.5 (Anthropic, 2024), GPT-4o-mini  (OpenAI, 2024), and Gemini 1.5  (Google, 2025) struggle with fine-grained code semantics. They often fail to reason about even single statements from real-world code—particularly arithmetic expressions and API calls—and perform poorly on tasks involving loop values and iteration counts. Basic chain-of-thought prompting offers limited benefit, and few-shot prompting yields only modest improvements. In-context learning is most effective when prompts define new concepts or include highly relevant examples. Interestingly, models can correlate natural language code semantics questions with certain code patterns. For instance, when the code contains assignments like p = q, models correctly respond to the prompt "do p and q at <line> alias the same memory address" even in zero-shot settings. Similarly, models reliably infer loop bounds in explicit cases such as for i in range(100):. Among the 14 models evaluated, Claude 3.5 consistently achieved the best performance. We also observe that Java and Python code are generally easier for models to reason about than C, and that input prediction (i.e., reverse semantic inference) remains among the most challenging tasks.
@@ 2 Benchmark Construction
Figure 2: CodeSense: data curation and benchmark creation overview
@@ 2.1 Defining a Spectrum of Code Reasoning Tasks
Task 1: Block-level code semantics (RQ1): To investigate whether a model understand a chuck of code, we give a block of statements. We give input and ask the models to predict the execution output; also we give output, we ask the models to predict the input. Input/output prediction of a function is a special case of probing block-level code semantics. In our evaluation, we sampled a block of statements from the entry of the functions and increased the sizes of blocks, including the entire function.
@@
Task 3-3: Branch property. Giving a function and its input as well as the location of a conditional branch in the function, we ask the model to predict what is the outcome of the branch. In our evaluation, we randomly selected a conditional branch in the function for prediction.
@@ 2.2 Collecting and Tracing Real-world Multi-Lingual Software Projects
It should be noted that although our repository are in the training dataset. However, the models are asked to predict dynamic information, which we generated by running test inputs over programs, using tools like fuzzers. Those values are not in the repos and have never been seen by the models. In fact, this is the advantage of our benchmark. The future benchmark designers can use our testing and tracing tools to run other inputs to freshly generate more data that are new to the models.
@@ 2.3 Data Filtering
We collected whole program traces, from which we curated unique functions based on their entry and exit points in the execution trace logs. We excluded functions that only contains comments, too lengthy to fit into the models’ context, and the functions which don’t have meaningful functionality in their body. For example, some functions only contain one statement like “return 0”, or “printf(“…”) or some functions are just a wrapper for another function which we have tested, such as "void myfunc(){ func();}". We obtained a total of 2125 Python, 876 C and 875 Java unique functions, with the sizes ranging from 3 to 516 lines of code. From these unique functions, we curated our task-specific datasets.
In real-world code, we face many complexities, e.g., the input of a function and the values of a variable can be complex types. As the first step of probing models to reason about fine-grained code semantics for real-world code, we focus on ground truth values of primitive data types in all tasks, including int, float, str, bool, list, pointer, double, dictionary, tuple, etc. In the evaluation, we show that even for values of primitive types, the models face challenging to predict them. We collected a total of 4483 samples from Python, C and Java, and constructed the ground truth for the above tasks and used for evaluation. See Table 2.
@@ 3 Evaluation
We evaluate 14 SOTA LLMs, 8 reasoning models and 6 non-reasoning models (Table  3), including open-source models (Llama, phi), close-source/API models (GPT-4.0 Mini, Claude 3.5 and Gemini 1.5) and distilled models (DeepSeek R1 series), with the model parameter sizes ranging from 7 to 14 billions. We utilized vLLM(v0.3.3) as our inference engine to run the models.
@@
We designed five different natural language prompt templates (see Appendix/data package), and ran them on a sampled dataset for each model. We observed that prompt templates are model-sensitive, but not task-sensitive. So we select a template for each model for all the tasks. We prompted the models to give a response inside specific tags (<ans> </ans>) and considered the response inside that tag to compare with the ground truth, as done in  Gu et al. (2024). For our evaluation metrics, we used accuracy (exact matching of the generated outputs of the models and the ground truth label).
@@ 3.2 Results for RQ2: Statement-level code semantics
We observe that arithmetic and API/calls are the most statement types, even for the best reasoning models like GPT4.0-mini and Claude 3. For APIs, we sampled frequently used third-party libraries, like os, sys, time and math installed by pip, but the models do not have knowledge about their execution semantics. We also experimented with adding the API definitions in the prompt, but it didn’t increase the performance significantly A.9. Models handle better for predicting Boolean values such as the output of a comparison statement, and also statements where constant is assigned to a variable. The models may understand the assignment operator "=" and have captured easy patterns like "a=3 indicates a has value 3 after executing this statement". We did not observe significant advantage of reasoning models over non-reasoning models for this task.
@@ 4 Related Work
Code Reasoning Benchmarks. CruxEval Gu et al. (2024) assesses LLMs’ performance on synthetic Python programs using the task of input-output prediction for a function. CruxEval-X Xu et al. (2024) extends this work to multilingual settings by translating synthetic Python programs in CruxEval to other languages using LLMs. REval Chen et al. (2024) evaluated branch prediction tasks using ClassEval and HumanEval Chen et al. (2021a). CodeMind Liu et al. (2024) proposed output prediction and code synthesis tasks on existing code benchmarks Chen et al. (2021b); Austin et al. (2021); Gu et al. (2024); Puri et al. (2021). They found that LLM reasoning capabilities deteriorate as program complexity increases Liu et al. (2024); Zhang et al. (2024a). Our benchmark CodeSense is the first that used real-world Python, C and Java code to evaluate LLMs’ reasoning capabilities. While most code reasoning benchmarks reason about function level execution semantics, we proposed and made ground truth available for a spectrum of fine-grained reasoning tasks regarding program behaviors within a function.
Other Code Application Benchmarks. SWE-Bench Jimenez et al. (2024) used the task of generating patches to resolve a given GitHub issues for real-world Python projects. SWE-PolyBench Rashid et al. (2025) extends this work to other programming languages. KGym Mathai et al. (2024) delivered a benchmark consisting of Linux kernel crash data and evaluated LLMs’ capabilities of resolving Linux kernel crashes. These benchmarks focus on task-specific performance rather than fine-grained code semantics understanding. There are also benchmarks for code generation Li et al. (2024); Zhang et al. (2024b); Yu et al. (2024); Chen et al. (2025); Du et al. (2023) and code completion Izadi et al. (2024); Ding et al. (2023b). However, most of these datasets—such as BigCodeBench Zhuo et al. (2024) and CodeBenchGen Xie et al. (2024) are restricted to a single language (primarily Python) and extracted from isolated competitive programming problems.
@@ A.1 Computation Resources and Inference Tools
  • Software Stack: vLLM  Kwon et al. [2023] for optimised transformer inference
@@ A.2 Limitations and Future Work
In this work, we consider exact matching between the model’s response and ground truth as correct. In future work, we would like to explore other metrics such as pass@k. However, we did explore models’ performance in computing abstract value prediction/approximation of code semantics.
@@
Additionally, we would like to expand our framework to support project tracing and task-specific benchmark datasets beyond the scope of three languages. This includes adding other languages with diverse domain’s like (e.g., functional, system language) to enable a more comprehensive evaluation of LLMs.
In future, it will be also interesting to explore more advanced prompting techniques and even fine-tune the models to further evaluate the models.

### v3

@@ Abstract
Understanding and reasoning about code semantics is essential for enhancing code LLMs’ abilities to solve real-world software engineering (SE) tasks. Although several code reasoning benchmarks exist, most rely on synthetic datasets or educational coding problems and focus on coarse-grained reasoning tasks such as input/output prediction, limiting their effectiveness in evaluating LLMs in practical SE contexts. To bridge this gap, we propose CodeSense , the first benchmark that makes available a spectrum of fine-grained code reasoning tasks concerned with the software engineering of real-world code. We collected Python, C and Java software projects from real-world repositories. We executed tests from these repositories, collected their execution traces, and constructed a ground truth dataset for fine-grained semantic reasoning tasks. We then performed comprehensive evaluations on state-of-the-art LLMs. Our results show a clear performance gap for the models to handle fine-grained reasoning tasks. Although prompting techniques such as chain-of-thought and in-context learning helped, the lack of code semantics in LLMs fundamentally limits models’ capabilities of code reasoning. Besides dataset, benchmark and evaluation, our work produced an execution tracing framework and tool set that make it easy to collect ground truth for fine-grained SE reasoning tasks, offering a strong basis for future benchmark construction and model post training. Our code and data are located at https://codesense-bench.github.io/.
@@ 1 Introduction
Semantic code reasoning—the capacity to understand and predict the behaviour of software—is a core requirement underpinning a wide range of complex software engineering (SE) tasks, including test input generation, vulnerability detection, fault localization, bug repair, refactoring, and functional verification. Unlike syntactic pattern matching, which may rely on token-level similarity or statistical regularities, semantic reasoning ("codesense") entails a deep, execution-oriented understanding of how software operates. Although code semantics can be expressed in many ways, in practice, developers engage in semantic reasoning through tasks like predicting a function’s input-output behavior, tracing variable values, analyzing control flow paths, identifying loop invariants, etc. This form of reasoning aligns with formal definitions from programming language theory—particularly operational semantics, which models step-by-step execution, and axiomatic semantics, which uses logical assertions to describe program properties. Such reasoning tasks also reflect the real-world demands placed on developers and provide a natural grounding for their day-to-day work.
@@
To this end, we propose CodeSense, a benchmark for fine-grained code semantic reasoning, constructed from real-world GitHub projects in Python, C, and Java (see Table 1). CodeSense  introduces a spectrum of reasoning tasks at statement, code-block, and function levels, targeting essential semantic properties frequently needed across SE activities. For example, predicting loop iteration counts is critical for input/output prediction, performance analysis, and detecting infinite loops (e.g., denial-of-service vulnerabilities). Branch condition prediction and reasoning about pointers in C code are important for test input generation and memory safety assurance. As illustrated in Figures 1(a) and 1(b), fine-grained semantic reasoning about arithmetic operations, control flow, and API semantics is foundational for a variety of SE applications. Prior work Ding et al. (2023a; 2024) has shown that incorporating semantic signals during training improves model performance on code generation, branch prediction, code clone detection, program repair and vulnerability detection tasks, motivating our design of CodeSense to comprehensively evaluate models’ capabilities for semantic reasoning.
@@
Table 1: Optimal design space of code reasoning benchmarks (○ denotes not support, ◐ denotes partial support, and ● denotes fully support)
Benchmark Real-World Projects Multi-lingual Function I/O Fine-Grained Reasoning Exec. Steps API Understanding Multi-File Context Realistic Project Structure
CruxEval (Gu et al., 2024)
CruxEval-X (Xu et al., 2024)
REval Chen et al. (2024)
CodeMind (Liu et al., 2024)
CoRe (Xie et al., 2025)
@@
To construct CodeSense , we collected 544 Python, 100 C, and 100 Java projects. To enable supervised evaluation, we develop a framework that automatically extracts ground-truth annotations for these tasks using static and dynamic program analysis. We build, execute and log run-time execution values and traces of the real-world projects. Our dataset includes a total of 2125 Python, 876 C, and 875 Java unique functions, based on which, we curated 4495 samples with their ground truth in our benchmark. Please see more details in Section 2.
Using our benchmark, we evaluated 14 state-of-the-art (SOTA) LLMs and investigated six research questions regarding the models’ code semantics reasoning capabilities. Previous work has shown that models did not perform well on code reasoning tasks such as input/output prediction Gu et al. (2024) and vulnerability detection Steenhoek et al. (2025); to understand why models fail and identify places for improvement, we investigate: RQ1: Does increasing code size make semantic reasoning more difficult? RQ2: Which types of program statements are easier or harder for models to reason about? RQ3: How do models perform on code properties critical for SE tasks, such as predicting pointer aliasing, loop iteration counts, and branch conditions? RQ4: How effective are different prompting strategies in improving semantic reasoning? RQ5: Can models reason approximately when exact values or semantics are hard to infer? RQ6: Do models perform better on some programming languages than others?
Our results reveal that current LLMs, including SOTA models like Claude 3.5  Anthropic (2024), GPT-4o-mini  OpenAI (2024), and Gemini 1.5  Google (2025) struggle with fine-grained code semantics. They often fail to reason about even single statements from real-world code—particularly arithmetic expressions and API calls—and perform poorly on tasks involving loop values and iteration counts. Basic chain-of-thought prompting offers limited benefit, and few-shot prompting yields only modest improvements. In-context learning is most effective when prompts define new concepts or include highly relevant examples. Interestingly, models can correlate natural language code semantics questions with certain code patterns. For instance, when the code contains assignments like p = q, models correctly respond to the prompt "do p and q at <line> alias the same memory address" even in zero-shot settings. Similarly, models reliably infer loop bounds in explicit cases such as for i in range(100):. Among the 14 models evaluated, Claude 3.5 consistently achieved the best performance. We also observe that Java and Python code are generally easier for models to reason about than C, and that input prediction (i.e., reverse semantic inference) remains among the most challenging tasks.
@@ 2.1 Defining a Spectrum of Code Reasoning Tasks
Figure 2: CodeSense: Data curation and benchmark creation overview.
Task 1: Block-level code semantics (RQ1): To investigate whether a model understand a chunck of code, we give a block of statements. We give input and ask the models to predict the execution output; also we give output, we ask the models to predict the input. Input/output prediction of a function is a special case of probing block-level code semantics. In our evaluation, we sampled a block of statements from the entry of the functions and increased the sizes of blocks, including the entire function.
@@
Task 3-3: Branch property. Given a function and its input as well as the location of a conditional branch in the function, we ask the model to predict what is the outcome of the branch. In our evaluation, we randomly selected a conditional branch in the function for prediction.
@@ 2.3 Data Filtering
We collected whole program traces, from which we curated unique functions based on their entry and exit points in the execution trace logs. We excluded functions that only contain comments, too lengthy to fit into the models’ context, and the functions which don’t have meaningful functionality in their body. For example, some functions only contain one statement like “return 0”, or “printf(“…”) or some functions are just a wrapper for another function which we have tested, such as "void myfunc(){ func();}". We obtained a total of 2125 Python, 876 C and 875 Java unique functions, with the sizes ranging from 3 to 516 lines of code. From these unique functions, we curated our task-specific datasets.
In real-world code, we face many complexities, e.g., the input of a function and the values of a variable can be complex types. As the first step of probing models to reason about fine-grained code semantics for real-world code, we focus on ground truth values of primitive data types in all tasks, including int, float, str, bool, list, pointer, double, dictionary, tuple, etc. In the evaluation, we show that even for values of primitive types, the models find it challenging to predict them. We collected a total of 4483 samples from Python, C and Java, and constructed the ground truth for the above tasks and used for evaluation. See Table 2.
@@ 3 Evaluation
We evaluate 14 SOTA LLMs, 8 reasoning models and 6 non-reasoning models (Table 3 for full names and short IDs used in figures), including open-source models (Llama, phi), close-source/API models (GPT-4.0 Mini, Claude 3.5 and Gemini 1.5) and distilled models (DeepSeek R1 series), with the model parameter sizes ranging from 7 to 14 billions. We utilized vLLM(v0.3.3) as our inference engine to run the models.
@@
We designed five different natural language prompt templates (see Appendix/data package), and ran them on a sampled dataset for each model. We observed that prompt templates are model-sensitive, but not task-sensitive. So we select a template for each model for all the tasks. We prompted the models to give a response inside specific tags (<ans> </ans>) and considered the response inside that tag to compare with the ground truth, as done in (Gu et al., 2024). For our evaluation metrics, we used accuracy (exact matching of the generated outputs of the models and the ground truth label).
@@ 3.2 Results for RQ2: Statement-level code semantics
We observe that arithmetic and API/calls are the most statement types, even for the best reasoning models like GPT4.0-mini and Claude 3. For APIs, we sampled frequently used third-party libraries, like os, sys, and math installed by pip, but the models do not have knowledge about their execution semantics. We also experimented with adding the API definitions in the prompt, but it didn’t increase the performance significantly A.9. Models handle better for predicting Boolean values such as the output of a comparison statement, and also statements where constant is assigned to a variable. The models may understand the assignment operator "=" and have captured easy patterns like "a=3 indicates a has value 3 after executing this statement". We did not observe significant advantage of reasoning models over non-reasoning models for this task.
@@ 4 Related Work
Code Reasoning Benchmarks. CruxEval (Gu et al., 2024) assesses LLMs’ performance on synthetic Python programs using the task of input-output prediction for a function. CruxEval-X (Xu et al., 2024) extends this work to multilingual settings by translating synthetic Python programs in CruxEval to other languages using LLMs. REval (Chen et al., 2024) evaluated branch prediction tasks using ClassEval and HumanEval (Chen et al., 2021a). CodeMind (Liu et al., 2024) proposed output prediction and code synthesis tasks on existing code benchmarks (Chen et al., 2021b; Austin et al., 2021; Gu et al., 2024; Puri et al., 2021). They found that LLM reasoning capabilities deteriorate as program complexity increases (Liu et al., 2024; Zhang et al., 2024a). Our benchmark CodeSense is the first that used real-world Python, C and Java code to evaluate LLMs’ reasoning capabilities. While most code reasoning benchmarks reason about function level execution semantics, we proposed and made ground truth available for a spectrum of fine-grained reasoning tasks regarding program behaviors within a function.
Other Code Application Benchmarks. SWE-Bench (Jimenez et al., 2024) used the task of generating patches to resolve a given GitHub issues for real-world Python projects. SWE-PolyBench (Rashid et al., 2025) extends this work to other programming languages. KGym (Mathai et al., 2024) delivered a benchmark consisting of Linux kernel crash data and evaluated LLMs’ capabilities of resolving Linux kernel crashes. These benchmarks focus on task-specific performance rather than fine-grained code semantics understanding. There are also benchmarks for code generation (Li et al., 2024; Zhang et al., 2024b; Yu et al., 2024; Chen et al., 2025; Du et al., 2023) and code completion (Izadi et al., 2024; Ding et al., 2023b). However, most of these datasets—such as BigCodeBench (Zhuo et al., 2024) and CodeBenchGen (Xie et al., 2024) are restricted to a single language (primarily Python) and extracted from isolated competitive programming problems.
@@ A.1 Computation Resources and Inference Tools
  • Software Stack: vLLM  [Kwon et al., 2023] for optimised transformer inference
@@ A.2 Limitations and Future Work
In this work, we consider exact matching between the model’s response and ground truth as correct. In future work, we would like to explore other metrics such as pass@k. However, we did explore models’ performance of computing abstract value prediction/approximation of code semantics.
@@
Additionally, we would like to expand our framework to support project tracing and task-specific benchmark datasets beyond the scope of three languages. This includes adding other languages with diverse domains like (e.g., functional, system language), to enable a more comprehensive evaluation of LLMs.
In future, it will be also be interesting to explore more advanced prompting techniques and even fine-tune the models to further evaluate them.
@@ A.9 API Definition Ablation Study
A.10 Comparison with Existing Benchmarks
We did a partial evaluation of the output prediction task on our evaluation framework on the three best-performing open-source models on CodeSense, and we present the results in Table 9. Our result shows that the models perform significantly worse in CodeSense than CruxEval.
Table 9: Output prediction accuracy comparison: CruxEval vs. CodeSense
Model CruxEval CodeSense Drop
DeepSeek-R1-Distill-Qwen-14B 0.75 0.37 0.38
Qwen2.5-14B 0.52 0.27 0.25
Qwen2.5-Coder-7B 0.50 0.30 0.20
A.11 Variance Analysis
We have run the statement prediction task, and the results in Table 10 across three runs demonstrate that the variance across multiple runs is minimal, and this doesn’t change our core findings.
Table 10: Variance analysis across three runs on statement prediction task
Model Run 1 Run 2 Run 3 Mean ± Std Dev
Qwen2.5-14B 44.4% 44.4% 43.3% 44.0% ± 0.6%
DeepSeek-R1-Distill-Qwen-14B 42.0% 41.8% 43.8% 42.5% ± 1.0%
Qwen2.5-Coder-7B 38.4% 38.4% 38.4% 38.3% ± 0.0%
DeepSeek-R1-Distill-Llama-8B 26.4% 25.7% 27.3% 26.5% ± 0.8%
A.12 Different Prompting techniques for Statement Prediction
Table 11 shows the difference between two prompting strategies: using in-context examples of the same statement type as the query versus examples of a different statement type.
Table 11: Statement Prediction Performance by Type
Model Same Type Statement Different Type Statement
Qwen2.5-14B-Instruct-1M 0.44 0.42
DeepSeek-R1-Distill-Qwen-14B 0.42 0.39
Qwen2.5-Coder-7B-Instruct 0.38 0.37
Llama-3.1-8B-Instruct 0.32 0.29
Phi-4-mini-instruct 0.30 0.25
A.13 Function Size Analysis
Table 12 shows how we categorise function difficulties based on lines of code.
Table 12: Function Size Categories
Category Length (Lines of Code)
Small $\text{length}\leq 9$
Medium $10<\text{length}\leq 19$
Large $\text{length}\geq 20$
A.14 Examples of Fine-Grained Insights
Codesense’s fine-grained task can uncover reasoning failures on code semantics that are not visible to the coarse-grained benchmarks. For example, consider the simple function from our benchmark
def _is_ascii(s):
    if isinstance(s, str):
        for c in s:
            if ord(c) > 255:
                return False
        return True
    return _supports_unicode(s)
_is_ascii(’ 123456789#’)
Question: “How many times will the loop on line 3 iterate?”
Ground Truth: 11 (The length of the input string s, which contains a leading space, digits 1–9, and the # character).
However, models such as Qwen2.5-Coder-7B incorrectly respond with 10. This error reveals that the model fails at a fundamental level: it cannot correctly reason about string iteration, specifically miscounting the characters in a simple string literal.